47 An engineers salary was $40,000 in 2004. The same engineers salary in 2011 is $75,000. If the companys salary policy di
Answer #1

Engineer's salary in 2004 = $40,000

Engineer's salary in 2011 = $75,000

Growth rate in the period 2004-12 = (Salary in 2011-Salry in 2004)/Salary in 2004 x 100 = ($75,000-$40,000)/$40,000 x 100 = $35,000/$40,000 x 100 = 87.5%

Let the average rate of inflation per year in the 8-year period be n%

(1+n%)8 - 1 = 87.5%

=> (1+n/100)8 = 87.5/100 + 1

=> (1+n)8 = 0.875 + 1

=> ((1+n)8)1/8 = 1.8751/8

=> 1+n = 1.08174

=> n = 0.0817 = 8.174%

Ans: 8.174%